# Einstein connection of nonsymmetric pseudo-Riemannian manifold

### v1

@@ Abstract
A. Einstein considered a nonsymmetric (0,2)-tensor $G=g+F$, where $g$ is a pseudo-Riemannian metric and $F\neq 0$ is skew-symmetric, and a linear connection $\nabla$ with torsion $T$ such that $(\nabla_{X}\,G)(Y,Z)=-G(T(X,Y),Z)$. M. Prvanović (1995) obtained the explicit form of the Einstein connection of an almost Hermitian manifold. In this paper, first, we present the result above in coordinate-free form, and then extend it to almost contact metric manifolds satisfying the so-called $f^{2}$-torsion condition. We then derive the Einstein connection of nonsymmetric pseudo-Riemannian, in particular, weak almost Hermitian manifolds $(M,f,g)$, satisfying the $f^{2}$-torsion condition, where $F(X,Y)=g(X,fY)$, give explicit formulas for the torsion in terms of $\nabla^{g}F$, $dF$ and a new (1,1)-tensor $\widetilde{Q}:=-f^{2}-{\rm Id}$, and show that in the almost Hermitian case, our main result reduces to the coordinate-free form of Prvanović’s solution. Finally, we describe special Einstein connections, i.e., the difference tensor has the property $K_{X}Y=-K_{Y}X$, and indicate the Gray-Hervella classes. Illustrative examples are given, including the construction of a weighted product.
@@ 1 Introduction
called an Einstein connection. The idea of a nonsymmetric basic tensor was revisited by J.W. Moffat [7], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [6] point to potential improvements and challenges within NGT. M.I. Wanas et al. [11] studied path equations of a particle in Einstein’s nonsymmetric geometry. Recent approaches to modified gravity often rely on differential geometry, including torsion and non-metricity, as a natural extension of General Relativity. Connections with totally skew-symmetric torsion, i.e., $T(X,Y,Z)$ is a 3-form, satisfying T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), are important due to the relations with mathematical physics, as they admit a coupling to spinor fields and lead to holonomy and rigidity phenomena. In this case, the dimension of the space of parallel spinors coincides with the number of preserved supersymmetries. Such connections arise naturally in supersymmetric string theories and non-linear $\sigma$-models. and references therein), as well as in gravitational models [4]. S. Ivanov and M. Lj. Zlatanović (see [5]) presented conditions for the existence and uniqueness of the Einstein connection with totally skew-symmetric torsion on a manifold $(M,G=g+F)$ and gave its explicit expression using an almost contact metric structure. In this case, the torsion is determined by $dF$: $T(X,Y,Z)=-\frac{1}{3}\,dF(X,Y,Z)$, and the difference tensor is given by
@@
Weak metric structures, developed by V. Rovenski et al., see [9], generalize the classical metric $f$-structures and are well suited for a study of $G=g+F$ in NGT with an arbitrary skew-symmetric tensor $F$. A weak almost Hermitian manifold $(M,f,g)$ is a Riemannian manifold $(M,g)$ of dimension $2m\geq 4$ endowed with non-singular skew-symmetric endomorphism $f$ and the 2-form $F$ such that $F(X,Y)=g(X,fY)$, see [12]. V. Rovenski and M. Zlatanović [12, 10] were the first to apply the weak $f$-structure, in particular, the weak almost contact structure, to NGT of totally skew-symmetric torsion with $F(X,Y)=g(X,fY)$.
@@
The paper has an Introduction and six sections. In Section 2 we review basics of NGT, give examples, prove several lemmas, and discuss the relation between torsion and the contorsion, $K$, of an Einstein connection emphasizing the role of $f^{2}$-torsion condition. The coordinate-free representation of M. Prvanović solution (Theorem 3.1) is given in Section 3. In Section 4 we deduce equations for Einstein connection of an almost contact metric manifold with the $f^{2}$-torsion condition (Theorem 4.2). In Section 5 we generalize the result of M. Prvanović and deduce equations for Einstein connection with the $f^{2}$-torsion condition of nonsymmetric pseudo-Riemannian manifold (Theorem 5.3), in particular, for weak almost Hermitian manifold (Corollary 5.5). Long expressions of Section 5 are given in Section 6.
@@ 2 Einstein’s nonsymmetric geometry
A linear connection $\nabla:\mathfrak{X}_{M}\times\mathfrak{X}_{M}\to\mathfrak{X}_{M}$ on a smooth manifold $M$ has the properties
$\nabla_{fX_{1}+X_{2}}Y=f\,\nabla_{X_{1}}Y+\nabla_{X_{2}}Y,\quad\nabla_{X}(fY_{1}+Y_{2})=f\,\nabla_{X}Y_{1}+X(f)\cdot Y_{1}+\nabla_{X}Y_{2}$
for any vector fields $X,Y,X_{1},X_{2},Y_{1},Y_{2}$ and a smooth function $f$ on $M$.
@@
Separating symmetric and skew-symmetric parts of (1.1), we express the covariant derivatives $\nabla g$ and $\nabla F$ in terms of the (0,3)-torsion tensor $T(X,Y,Z):=g(T(X,Y),Z)$:
$(\nabla_{X}\,g)(Y,Z)$ $=\tfrac{1}{2}\big[T(Z,X,\,Y+fY)-T(X,Y,\,Z+fZ)\big],$ (2.2)
$(\nabla_{Z}\,F)(X,Y)$ $=-\tfrac{1}{2}\big[T(Z,X,\,Y+fY)+T(Y,Z,\,X+fX)\big].$ (2.3)
Recall the co-boundary formula for a 2-form $F$, for example, [1]:
@@
Note that the equality (2) yields
@@
The Einstein connection $\nabla$ is represented in [5] using the torsion $T$ as
@@
By (2.10), the difference (0,3)-tensor $K(X,Y,Z):=g(K_{X}Y,Z)$ and torsion tensor of an Einstein connection $\nabla$ are expressed linearly in terms of each other:
@@
Note that, see (2.1),
@@
Let $(M,f,g)$ be a weak nearly Kähler manifold, i.e., considered as a generalized Riemannian manifold $(M,G=g+F)$. Suppose that $\nabla$ is an Einstein connection on $(M,f,g)$ with totally skew-symmetric torsion. Then $\nabla$ satisfies the $f$-torsion condition:
@@
or, equivalently, $T(fX,Y,Z)=T(X,fY,Z)=T(X,Y,fZ)$, and torsion of $\nabla$ is determined by $dF$ as $T(X,Y,Z)=-\tfrac{1}{3}\,dF(X,Y,Z)$, see Example 3.4 in [12]. In this case, (1.2) reduces to $K_{X}Y=\tfrac{1}{2}\,T(X,Y)$, and $\nabla$ is recovered by the equality $\nabla_{X}Y=\nabla^{g}_{X}Y+\tfrac{1}{2}\,T(X,Y)$.
@@
If a manifold $(M,G=g+F)$ admits an Einstein connection $\nabla$ with totally skew-symmetric torsion, and the $f$-torsion condition (2.17) holds, then (2.19) is true. Indeed, using (2.17) and the totally skew-symmetric torsion property in the LHS of (2.18) gives zero.
The following skew-symmetry, obtained using (2.12):
$K_{X}Y=-K_{Y}X\qquad\text{or equivalently}\qquad K_{X}Y=\tfrac{1}{2}\,T(X,Y),$ (2.20)
characterizes special Einstein connections, see [8]. This condition expresses the skew–symmetry of $K$ only in the first two arguments, and doesn’t imply that the total skew–symmetry of $K(X,Y,Z)$, which requires, in addition, condition (2.18). In this case, torsion tensor becomes totally skew–symmetric and satisfies the $f$-torsion condition, see Lemma 2.5.
@@
If $K$ is totally skew-symmetric, then $T$ is totally skew-symmetric, and the $f$-torsion condition (2.17) is true.
@@
Let (2.18) be true. Then we get the equality
@@
therefore, $K$ is totally skew-symmetric if and only if $T$ is totally skew-symmetric. Moreover, the total skew-symmetry of $T$ implies the $f$-torsion condition (2.17).
@@
The Nijenhuis tensor $N_{f}$ is defined as
$N_{f}(X,Y)=[fX,fY]+f^{2}[X,Y]-f[fX,Y]-f[X,fY],$
and can be written in terms of $\nabla^{g}f$ as
$N_{f}(X,Y)=(f\nabla^{g}_{Y}f-\nabla^{g}_{fY}f)X-(f\nabla^{g}_{X}f-\nabla^{g}_{fX}f)Y.$
Using (2.18) in (2.3), we obtain
$(\nabla_{X}F)(Y,Z)=-T(X,Y,Z)-T(X,Y,fZ).$ (2.24)
By (2.24), $\nabla F=0$ (when $F$ is non-degenerate) implies $T=0$.
@@
3 Explicit form of the Einstein connection of almost Hermitian manifold
In this section, we represent the solution (an explicit form of the Einstein connection of an almost Hermitian manifold) of M. Prvanović [8] in the coordinate-free form.
Theorem 3.1 (see Theorem 1 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a generalized Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$ Then an Einstein connection $\nabla$ on $M$ is given by
$-(\nabla^{g}_{Y}F)(Z,X)-(\nabla^{g}_{Z}F)(X,Y).$ (3.1)
Proof.
We calculate following the steps of [8]:
$(\nabla^{g}_{X}F)(JY,JZ)$ $=-(\nabla^{g}_{X}F)(Y,Z),$ (3.2)
$(\nabla^{g}_{X}F)(JY,Z)$ $=(\nabla^{g}_{X}F)(Y,JZ),$
and, see (2.10),
$g(\nabla_{X}Y,Z)$ $=g(\nabla^{g}_{X}Y,Z)+\tfrac{1}{2}\big\{T(X,Y,Z)-T(Z,X,JY)+T(Y,Z,JX)\big\}.$
Using $G=g+F$ and (1.1), we obtain, see (2.7):
$2\,(\nabla^{g}_{X}F)(Y,Z)=$ $\,-T(Z,X,Y)-T(X,Y,Z)-T(JZ,X,JY)$
$-T(X,JY,JZ)+T(Y,JZ,JX)+T(JY,Z,JX).$ (3.3)
Transvect (3.3):
$(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{X}F)(JY,JZ)=T(Y,JZ,JX)+T(JY,Z,JX).$
Substituting the formula above into (3.3) and using (3.2), we have
$T(X,Y,Z)+T(Z,X,Y)+T(X,JY,JZ)+T(JZ,X,JY)=0.$ (3.4)
Thus (3.3) reduces to
$2\,(\nabla^{g}_{X}F)(Y,Z)=T(Y,JZ,JX)+T(JY,Z,JX).$ (3.5)
The cyclic sum of (3.5) is
$\quad(\nabla^{g}_{X}F)(Y,Z)+(\nabla^{g}_{Y}F)(Z,X)+(\nabla^{g}_{Z}F)(X,Y)$
$=-T(X,Y,Z)-T(Y,Z,X)-T(Z,X,Y).$ (3.6)
The relation (3) with $Y\to JY$ and $Z\to JZ$, in view of (3.4), implies:
$\quad(\nabla^{g}_{X}F)(JY,JZ)+(\nabla^{g}_{JY}F)(JZ,X)+(\nabla^{g}_{JZ}F)(X,JY)$
$=T(X,Y,Z)+T(Z,X,Y)-T(JY,JZ,X).$ (3.7)
Transvecting (3.5), we get the following:
$T(JY,JZ,X)=T(Y,Z,X)-2\,(\nabla^{g}_{JX}F)(JY,Z).$ (3.8)
Applying (3.8) to (3.7), gives
$\quad(\nabla^{g}_{X}F)(JY,JZ)+(\nabla^{g}_{JY}F)(JZ,X)+(\nabla^{g}_{JZ}F)(X,JY)$
$=T(X,Y,Z)+T(Z,X,Y)-T(Y,Z,X)+2(\nabla^{g}_{JX}F)(JY,Z).$
The relation above together with (3) and (3.2) gives (3.1). ∎
Corollary 3.2 (see Theorem 2 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a generalized Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$ Then a special Einstein connection $\nabla$ on $M$ is given by
$T(X,Y,Z)=\tfrac{1}{2}\big\{(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{JZ}F)(JX,Y)-(\nabla^{g}_{JY}F)(JX,Z)\big\}.$ (3.9)
Consequently, by (3.9), the special Einstein connection $\nabla$ is determined by
$g(\nabla_{X}Y,Z)=g(\nabla^{g}_{X}Y,Z)+\tfrac{1}{2}\big\{(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{JZ}F)(JX,Y)-(\nabla^{g}_{JY}F)(JX,Z)\big\}.$
Proof.
Equivalently,
$(\nabla^{g}_{X}F)(JY,Z)+(\nabla^{g}_{JY}F)(X,Z)+(\nabla^{g}_{Y}F)(JX,Z)+(\nabla^{g}_{JX}F)(Y,Z)=0.$
Thus, (3.1) reduces to (3.9), defining the torsion of the special Einstein connection. ∎
Remark $3.$
${\mathcal{W}}_{1},\ {\mathcal{W}}_{3},\ {\mathcal{W}}_{4},\ {\mathcal{W}}_{3}\oplus{\mathcal{W}}_{4},\ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{3},\ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{4},\ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{3}\oplus{\mathcal{W}}_{4}.$
4 Explicit form of the Einstein connection of almost contact metric manifold
Here, we deduce equations for Einstein connection for an almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$, developing the approach (of Section 3) for almost Hermitian manifolds.
We have, see, for example, [1],
$f^{2}$ $=-{\rm Id}+\eta\otimes\xi,$
$g(fX,fY)$ $=g(X,Y)-\eta(X)\eta(Y).$ (4.1)
Differentiating by $X$ the identity (4) yields
$\nabla^{g}_{X}f^{2}=\nabla^{g}_{X}(-{\rm Id}+\eta\otimes\xi)=\nabla^{g}_{X}(\eta\otimes\xi).$
Using the Leibniz rule, we have
$\nabla^{g}_{X}(\eta\otimes\xi)$ $=(\nabla^{g}_{X}\eta)\otimes\xi+\eta\otimes\nabla^{g}_{X}\xi,$
$\nabla^{g}_{X}f^{2}$ $=(\nabla^{g}_{X}f)\,f+f\,\nabla^{g}_{X}f.$
Applying the equality above to $Y$ yields
@@ 4 Explicit form of the Einstein connection of almost contact metric manifold
The following two conditions are true for all vector fields $X,Y,Z$ on an almost contact metric manifold:
@@
Since $\nabla^{g}g=0$ and $F(Y,Z)=g(Y,fZ)$, we have $(\nabla^{g}_{X}F)(Y,Z)=g\big(Y,(\nabla^{g}_{X}f)Z\big)$. Hence
@@
Therefore
$(\nabla^{g}_{X}F)(fY,fZ)$ $=-g\big(Y,(\nabla^{g}_{X}f)Z\big)+\eta(Y)\eta\big((\nabla^{g}_{X}f)Z\big)+\;\eta(Z)\,g(fY,\nabla^{g}_{X}\xi).$
Finally, we use
@@
Substituting this in equation above yields (4.3). Using $Z\to fZ$ in (4.3), we get (4.4). ∎
@@
Let $\nabla$ be an Einstein connection satisfying the $f^{2}$-torsion condition (2.27) on an almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$, considered as an NGT space $(M,G=g+F,\nabla)$. Then
@@
In particular, the torsion $T$ is horizontal, and for $X,Y,Z\perp\xi$ it is given by the same formula as for the almost Hermitian case (3.1):
@@
We write $Z=f^{2}Z^{\prime}+\eta(Z)\xi$, where $Z^{\prime}\perp\xi$, and by $f^{2}$-torsion condition (2.27), we get
@@
Using (4.7) and the $f^{2}$-torsion condition (2.27), we obtain
$T(X,Z,\xi)$ $=T(X,f^{2}Z^{\prime},\xi)+\eta(Z)\,T(X,\xi,\xi)=T(X,f^{2}Z^{\prime},\xi)=T(f^{2}X,Z^{\prime},\xi)=0,$
$T(\xi,Z,X)$ $=T(\xi,f^{2}Z^{\prime},X)+\eta(Z)\,T(\xi,\xi,X)=T(\xi,f^{2}Z^{\prime},X)=T(f^{2}\xi,Z^{\prime},X)=0.$
@@
The rest of proof follows Section 3, using (4.8) and (4.9). All equations from (3.4)-(3.8) are valid after replacing $J$ by $f$. Finally, since $g(\xi,\xi)=1$, we obtain
$(\nabla^{g}_{X}\eta)(Z)=g(\nabla^{g}_{X}\xi,Z)=-(\nabla^{g}F)(X,fZ,\xi)=0,$
hence $\nabla^{g}\,\xi=0$. ∎
@@
Although the $f^{2}$-torsion condition is trivial for almost Hermitian manifolds, it is strong enough for almost contact metric manifolds. In Theorem 4.2, our manifold is locally the direct product of a real line and an almost Hermitian manifold $(\bar{M}^{2n},J,\bar{g})$, and $\nabla$, being an Einstein connection on the $2n$-dimensional factor, is given by (3.1), where $J$ is replaced by $f$. Consequently, the 1-form $\eta$ is closed, $d\eta=0$. The distribution $\ker\eta$ is involutive; thus, it is Frobenius integrable. The integral submanifold $\bar{M}^{2n}$ is almost Hermitian.

### v2

# Einstein connection of nonsymmetric Riemannian manifold

@@ Abstract
A. Einstein considered a nonsymmetric (0,2)-tensor $G=g+F$, where $g$ is a pseudo-Riemannian metric and $F\neq 0$ is skew-symmetric, and a linear connection $\nabla$ with torsion $T$ such that $(\nabla_{X}\,G)(Y,Z)=-G(T(X,Y),Z)$. M. Prvanović (1995) obtained the explicit form of the Einstein connection of an almost Hermitian manifold. In this paper, first, we present the result above in coordinate-free form, and then extend it to almost contact metric manifolds satisfying the $f^{2}$-torsion condition $T(f^{2}X,Y)=T(X,f^{2}Y)=f^{2}T(X,Y)$, where $F(X,Y)=g(X,fY)$. We then derive the Einstein connection of nonsymmetric pseudo-Riemannian, in particular, weak almost Hermitian manifolds $(M,f,g)$, satisfying the $f^{2}$-torsion condition, give explicit formulas for the torsion in terms of $\nabla^{g}F$, $dF$ and a new (1,1)-tensor $\widetilde{Q}=-f^{2}-{I}$, and show that in the almost Hermitian case, our results reduce to the coordinate-free form of Prvanović’s solution. Finally, we discuss special Einstein connections, and indicate the Gray-Hervella classes. Illustrative examples are given, including the construction of a weighted product.
@@ 1 Introduction
called an Einstein connection. The idea of a nonsymmetric basic tensor was revisited by J.W. Moffat [7], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [6] point to potential improvements and challenges within NGT. M.I. Wanas et al. [11] studied path equations of a particle in Einstein’s nonsymmetric geometry. Recent approaches to modified gravity often rely on differential geometry, including torsion and non-metricity, as a natural extension of General Relativity. Connections with totally skew-symmetric torsion, i.e., $T(X,Y,Z)$ is a 3-form, satisfying T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), are important due to the relations with mathematical physics, as they admit a coupling to spinor fields and lead to holonomy and rigidity phenomena. In this case, the dimension of the space of parallel spinors coincides with the number of preserved supersymmetries. Such connections arise naturally in supersymmetric string theories and non-linear $\sigma$-models, as well as in gravitational models [4]. S. Ivanov and M. Lj. Zlatanović (see [5]) presented conditions for the existence and uniqueness of the Einstein connection with totally skew-symmetric torsion on a manifold $(M,G=g+F)$ and gave its explicit expression using an almost contact metric structure. In this case, the torsion is determined by $dF$: $T(X,Y,Z)=-\frac{1}{3}\,dF(X,Y,Z)$, and the difference tensor is given by
@@
Weak metric structures, developed by V. Rovenski et. al, see [9], generalize the classical metric $f$-structures and are well suited for a study of $G=g+F$ in NGT with an arbitrary skew-symmetric tensor $F$. A weak almost Hermitian manifold $(M,f,g)$ is a Riemannian manifold $(M,g)$ of dimension $2m\geq 4$ endowed with non-singular skew-symmetric endomorphism $f$ and the 2-form $F$ such that $F(X,Y)=g(X,fY)$, see [12]. V. Rovenski and M. Zlatanović [12, 10] were the first to apply the weak $f$-structure, in particular, the weak almost contact structure, to NGT of totally skew-symmetric torsion with $F(X,Y)=g(X,fY)$.
@@
The paper has an Introduction and six sections. In Section 2 we review basics of NGT, give examples, prove several lemmas, and discuss the relation between torsion and the contorsion, $K$, of an Einstein connection emphasizing the role of $f^{2}$-torsion condition. In Section 3 we deduce equations for Einstein connection of an almost contact metric manifold with the $f^{2}$-torsion condition (Theorem 3.3). In Section 4 we generalize the result of M. Prvanović and deduce equations for an Einstein connection with the $f^{2}$-torsion condition of nonsymmetric pseudo-Riemannian manifold (Theorem 4.3), in particular, for weak almost Hermitian manifold (Corollary 4.5). Long expressions of Section 4 are given in Section 5.
@@ 2 Einstein’s nonsymmetric geometry
Separating symmetric and skew-symmetric parts of (1.1), we express the covariant derivatives $\nabla g$ and $\nabla F$ in terms of the (0,3)-torsion tensor $T(X,Y,Z)=g(T(X,Y),Z)$:
$2\,(\nabla_{X}\,g)(Y,Z)$ $=T(Z,X,\,Y+fY)-T(X,Y,\,Z+fZ),$ (2.2)
$2\,(\nabla_{Z}\,F)(X,Y)$ $=-T(Z,X,\,Y+fY)-T(Y,Z,\,X+fX).$ (2.3)
Recall the co-boundary formula for a 2-form $F$ (without the coefficient 3, unlike [1]):
@@
The equality (2) yields
@@
The Einstein connection $\nabla$ is represented in [5, Eq. (3.7)] using the torsion $T$ as
@@
By (2.10), the difference (0,3)-tensor $K(X,Y,Z)=g(K_{X}Y,Z)$ and the torsion tensor of an Einstein connection $\nabla$ are expressed linearly in terms of each other:
@@
Using various presentations of tensors and (2.1), we obtain
@@
Let $(M,f,g)$ be a weak nearly Kähler manifold, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$ with $F(X,Y)=g(X,fY)$. If $\nabla$ is an Einstein connection on $M$ with totally skew-symmetric torsion $T$, then the $f$-torsion condition holds:
@@
and the torsion of $\nabla$ is determined by $T(X,Y,Z)=-\tfrac{1}{3}\,dF(X,Y,Z)$, see Example 3.4 in [12]. In this case, (1.2) reduces to $K_{X}Y=\tfrac{1}{2}\,T(X,Y)$, and $\nabla_{X}Y=\nabla^{g}_{X}Y+\tfrac{1}{2}\,T(X,Y)$ holds.
@@
If a nonsymmetric Riemannian manifold $(M,G=g+F)$ admits an Einstein connection $\nabla$ with totally skew-symmetric torsion, and (2.17) holds, then (2.19) is true. Indeed, using (2.17) and the totally skew-symmetric torsion property in the LHS of (2.18) gives zero.
The following skew-symmetry of the difference tensor $K$ is obtained using (2.12):
$K_{X}Y=-K_{Y}X,\quad\text{or, equivalently,}\quad K_{X}Y=\tfrac{1}{2}\,T(X,Y),$ (2.20)
and characterizes special Einstein connections, see [8]. This condition expresses the skew–symmetry of $K$ only in the first two arguments, and doesn’t imply the total skew–symmetry of $K(X,Y,Z)$, which requires, in addition, condition (2.18). In this case, torsion tensor becomes totally skew–symmetric and satisfies the $f$-torsion condition, see Lemma 2.5.
@@
If the tensor $K$ is totally skew-symmetric, then the torsion tensor $T$ is totally skew-symmetric, and the $f$-torsion condition (2.17) is true.
@@
Let the condition (2.18) be true. Then we get the equality
@@
therefore, the tensor $K$ is totally skew-symmetric if and only if the torsion $T$ is totally skew-symmetric. Moreover, the total skew-symmetry of $T$ implies the condition (2.17).
@@
We represent the solution (an explicit form of the Einstein connection of an almost Hermitian manifold) of M. Prvanović [8] in the coordinate-free form.
Theorem 2.8 (see Theorem 1 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$ Then an Einstein connection $\nabla$ on $M$ is given by
$-(\nabla^{g}_{Y}F)(Z,X)-(\nabla^{g}_{Z}F)(X,Y).$ (2.26)
Corollary 2.9 (see Theorem 2 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a generalized Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$. Then the torsion of a special Einstein connection $\nabla$ on $M$, see (2.20), is given by
$2\,T(X,Y,Z)=(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{JZ}F)(JX,Y)-(\nabla^{g}_{JY}F)(JX,Z).$
${\mathcal{W}}_{1},\ \ {\mathcal{W}}_{3},\ \ {\mathcal{W}}_{4},\ \ {\mathcal{W}}_{3}\oplus{\mathcal{W}}_{4},\ \ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{3},\ \ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{4},\ \ {\mathcal{W}}_{1}\oplus{\mathcal{W}}_{3}\oplus{\mathcal{W}}_{4};$
3 Einstein connection of an almost contact metric manifold
Contact Riemannian geometry is a significant field in both pure mathematics and theoretical physics, for example, [1]. Developing the approach of [8], we deduce equations for Einstein connection of an almost contact metric manifold $(M^{2n+1},f,\xi,\eta,g)$. Here, $g$ is a Riemannian metric, $f$ is a skew-symmetric (1,1)-tensor of rank 2n2\,n, $\xi$ is a unit (Reeb) vector field, and $\eta$ is a 1-form such that $\eta(\xi)=1$. These tensors satisfy the relations
$f^{2}=-{I}+\eta\otimes\xi,\quad g(fX,fY)=g(X,Y)-\eta(X)\eta(Y),$ (3.1)
where I{I} is the identity map. Differentiating by $X$ the first identity of (3.1), yields
$\nabla^{g}_{X}f^{2}=\nabla^{g}_{X}(-{I}+\eta\otimes\xi)=\nabla^{g}_{X}(\eta\otimes\xi).$
Using the Leibniz rule, we have
$\nabla^{g}_{X}(\eta\otimes\xi)=(\nabla^{g}_{X}\eta)\otimes\xi+\eta\otimes\nabla^{g}_{X}\xi,\qquad\nabla^{g}_{X}f^{2}=(\nabla^{g}_{X}f)\,f+f\,\nabla^{g}_{X}f.$
Applying the equalities above to $Y$, yields
@@ 3 Einstein connection of an almost contact metric manifold
The following conditions are true for all vector fields $X,Y,Z$ on an almost contact metric manifold:
@@
Using $\nabla^{g}g=0$ and $F(Y,Z)=g(Y,fZ)$, we have $(\nabla^{g}_{X}F)(Y,Z)=g\big(Y,(\nabla^{g}_{X}f)Z\big)$. Hence
@@
Therefore,
$(\nabla^{g}_{X}F)(fY,fZ)=-g\big(Y,\nabla^{g}_{X}f)Z\big)+\eta(Y)\eta\big((\nabla^{g}_{X}f)Z\big)+\eta(Z)\,g(fY,\nabla^{g}_{X}\xi).$ (3.6)
Finally, substituting the equalities
@@
in (3.6), yields (3.4). Taking $Z\to fZ$ in (3.4), we get (3.5). ∎
@@
Let $(M^{2n+1},f,\xi,\eta,g)$ be an almost contact metric manifold, considered as an NGT space $(M,G=g+F,\nabla)$ with $F(X,Y)=g(X,fY)$. If the Einstein connection $\nabla$ satisfies the $f^{2}$-torsion condition (3.3), then
@@
In particular, the torsion $T$ is horizontal, and for $X,Y,Z\perp\xi$ it is given by the same formula (2.8) as for the almost Hermitian case:
@@
We write $Z=f^{2}Z^{\prime}+\eta(Z)\xi$, where $Z^{\prime}\perp\xi$, and by (3.3), we get
@@
Using the $f^{2}$-torsion condition (3.3) again and (3.9), we obtain
$T(X,Z,\xi)$ $=T(X,f^{2}Z^{\prime},\xi)+\eta(Z)\,T(X,\xi,\xi)=T(f^{2}X,Z^{\prime},\xi)=0,$
$T(\xi,Z,X)$ $=T(\xi,f^{2}Z^{\prime},X)+\eta(Z)\,T(\xi,\xi,X)=T(f^{2}\xi,Z^{\prime},X)=0.$
@@
We calculate, following the approach of [8]:
$(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z),\qquad(\nabla^{g}_{X}F)(fY,Z)=(\nabla^{g}_{X}F)(Y,fZ).$ (3.14)
From (1.1) with $G=g+F$ we obtain
$2\,(\nabla^{g}_{X}F)(Y,Z)=$ $\,-T(Z,X,Y)-T(X,Y,Z)-T(X,fY,fZ)$
$-T(fZ,X,fY)+T(Y,fZ,fX)+T(fY,Z,fX).$ (3.15)
Using (3.15), we get
$(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{X}F)(fY,fZ)=T(Y,fZ,fX)+T(fY,Z,fX).$
Substituting this into (3.15) and using (3.14), we obtain
$T(X,Y,Z)+T(Z,X,Y)+T(X,fY,fZ)+T(fZ,X,fY)=0.$
Thus (3.15) reduces to
$2\,(\nabla^{g}_{X}F)(Y,Z)=T(Y,fZ,fX)+T(fY,Z,fX).$ (3.16)
The cyclic sum of (3.16) is
$(\nabla^{g}_{X}F)(Y,Z)+(\nabla^{g}_{Y}F)(Z,X)+(\nabla^{g}_{Z}F)(X,Y)$
$=-T(X,Y,Z)-T(Y,Z,X)-T(Z,X,Y).$ (3.17)
Replacing $Y\to fY$ and $Z\to fZ$ in (3) gives
$(\nabla^{g}_{X}F)(fY,fZ)+(\nabla^{g}_{fY}F)(fZ,X)+(\nabla^{g}_{fZ}F)(X,fY)$
$=T(X,Y,Z)+T(Z,X,Y)-T(fY,fZ,X).$ (3.18)
Replacing $X\to fX$ and $Y\to fY$ in (3.16), we obtain
$T(fY,fZ,X)=T(Y,Z,X)-2\,(\nabla^{g}_{fX}F)(fY,Z).$ (3.19)
Substituting (3.19) into (3), gives the formula (3.3). ∎
@@
Although the $f^{2}$-torsion condition becomes trivial for almost Hermitian manifolds, it is strong enough for almost contact metric manifolds. In Theorem 3.3, our manifold is locally the direct product of $\mathbb{R}$ and an almost Hermitian manifold $(\bar{M}^{2n},J,\bar{g})$, and $\nabla$, being an Einstein connection on the $2n$-dimensional factor, is given by (2.8), where $J$ is replaced by $f$. Consequently, the 1-form $\eta$ is closed, $d\eta=0$. The distribution $\ker\eta$ is involutive; thus, it is Frobenius integrable. The integral submanifold $\bar{M}^{2n}$ is almost Hermitian.